OpenStudy (anonymous):

Can you please help me with a little algebra? I just need to sprinkle some solving powder on it to get started...

OpenStudy (anonymous):

\[y = (x-1)/(2x+3)\]

OpenStudy (anonymous):

I am attempting to solve for x.

OpenStudy (anonymous):

But I find myself going in a circle instead of isolating x. Can you give me a hint?

OpenStudy (anonymous):

Multiply both sides by 2x+3, you should get (2x+3)y=x-1 2xy+3y=x-1 2xy+3y+1=x 2xy-x=-3y-1 x(2y-1)=-3y-1 x=(-3y-1)/(2y-1)

OpenStudy (anonymous):

Try factoring out - x instead, which will make for a positive numerator, and I like that better. Thanks for helping me get unstuck.

OpenStudy (anonymous):

Your welcome, how is that calc group doing?

OpenStudy (anonymous):

I never connected with anybody at my ideal time. I did finish my Calc 1 class locally, and just found out I got an A on the final.

OpenStudy (anonymous):

Thank you for remembering.

OpenStudy (anonymous):

Nice job:)

OpenStudy (anonymous):

I don't think I could have gotten an A at MIT, though. I am working through the OCW class on my own in prep for Calc 2 next semester. The teaching is better, the recitations are more logically connected to the future of calculus work coming in the class, and the problems are clearly better structured to teach something specific while doing the homework as a side benefit to practicing a particular kind of problem. That makes the OCW homework feel much much more rewarding.